# On RoCK blocks of double covers of symmetric and alternating groups and the refined Broué conjecture

### v1

@@ 1 Introduction
Let $0<d<p$ and $\rho$ be a $d$-Rouquier $\bar{p}$-core. Suppose that $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $\widetilde{S}_{n}$ is $\widetilde{S}_{n}^{-}$. Then $B^{\rho,d}_{R_{p}}$ and $(B^{\rho,d}_{R_{p}})_{\bar{0}}$ are splendidly Rickard equivalent to $b^{\rho,d}_{R_{p}}$ and $(b^{\rho,d}_{R_{p}})_{\bar{0}}$ respectively. In particular, Conjecture 1.1 holds for RoCK blocks of $\widetilde{S}_{n}^{+}$ and $\widetilde{A}_{n}$, and if $p\equiv 1~{\rm mod}~4$, Conjecture 1.1 also holds for RoCK blocks of $\widetilde{S}_{n}^{-}$.
@@
Remark 1.10.
For the group $\widetilde{S}_{n}^{-}$, if $p\equiv 3~{\rm mod}~4$, Theorems 1.2, 1.4, 1.6 and 1.7 do not descend to $R_{p}$, but descend to the subring of $R$ generated by $R_{p}$ and a primitive $4$-th root of unity $\sqrt{-1}$. The reason is that by (5.2) below we have an isomorphism of superalgebras
$R\widetilde{S}_{n}^{-}e_{z}\cong\mathcal{T}_{n}^{R},~t_{i}e_{z}\mapsto(-1)^{i}\sqrt{-1}\tau_{i},\ ze_{z}\mapsto-1,$
but the element $\sqrt{-1}\in R$ is not in the subring $R_{p}$.
Using [1, Theorem 1.4] and [14, Theorem 1], we obtain a corollary of Theorem 1.9:
@@
Refinements of the Alperin–McKay conjecture by Isaacs–Navarro, Navarro, Turull and Boltje (see [1, (1.5), (1.6), (1.7), (1.8)]) and the blockwise Navarro/Galois Alperin weight conjecture (see [14, Conjecture 2]) hold for RoCK blocks of $\widetilde{S}_{n}^{+}$ and $\widetilde{A}_{n}$, and if $p\equiv 1~{\rm mod}~4$, these conjectures also hold for RoCK blocks of $\widetilde{S}_{n}^{-}$.
The blockwise Navarro Alperin weight conjecture has been proved for all blocks of $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$ in [8] using different method.
Let $\mathbb{P}$ be the set of prime numbers, and $\mathbb{P}_{1}:=\{p\in\mathbb{P}\mid p\equiv 1~{\rm mod}~4\}$. Now consider the following three families of group-prime pairs: $\Omega_{1}:=\{(\widetilde{S}_{n}^{+},p)\mid n\in\mathbb{N},~p\in\mathbb{P}\}$, $\Omega_{2}:=\{(\widetilde{A}_{n},p)\mid n\in\mathbb{N},~p\in\mathbb{P}\}$, $\Omega_{3}:=\{(\widetilde{S}_{n}^{-},p)\mid n\in\mathbb{N},~p\in\mathbb{P}_{1}\}$. For a partition $\lambda$ we denote by $h(\lambda)$ the number of nonzero parts of $\lambda$. Now by the results (Theorems 1.3, 1.5, 1.8) in this paper, to prove Conjecture 1.1 for all spin blocks of these three families, it suffices to prove the following refined version of the Kessar–Schaps conjecture [20]:
@@
Arbitrary spin blocks $B^{\rho,d}_{R_{p}}$ and $B^{\rho^{\prime},d^{\prime}}_{R_{p}}$ of symmetric groups (or spin blocks $(B^{\rho,d}_{R_{p}})_{\bar{0}}$ and $(B^{\rho^{\prime},d^{\prime}}_{R_{p}})_{\bar{0}}$ of alternating groups) are splendidly Rickard equivalent if and only if $d=d^{\prime}$ and $|\rho|-h(\rho)\equiv|\rho^{\prime}|-h(\rho^{\prime})~({\rm mod}~2)$.
@@
This paper is organised as follows: In Section 2 we review some relevant results on superalgebras and supermodules which are necessary for later use. In Section 3, we prove general descent results for supermodules and superequivalences. In Section 4, we formulate convenient reference for results on graded Rickard equivalences and splendid complexes. In Section 5 we review some block theory of $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$. Base on [21], we prove in Section 6 the structure results for Brauer correspondents of spin blocks over $\mathbb{F}_{p}$. In Section 7, we analyse the Morita superequivalence bimodule for an RoCK block constructed by Kleshchev and Livesey [22], and then realise this bisupermodule over $\mathbb{Z}_{p}$; we prove the refined Broué conjecture for RoCK spin blocks in the end of Section 7.
@@ 2 Preliminaries
Generalities. Throughout the rest of this paper $p$ is an odd prime, $\mathcal{O}$ is a complete discrete valuation ring with residue field $k$ of characteristic $p$ and quotient field $K$ of characteristic 0. Such a triple $(K,\mathcal{O},k)$ is called a $p$-modular system. (Note that we didn’t assume $(K,\mathcal{O},k)$ to be large enough.) An extension of the $p$-modular system $(K,\mathcal{O},k)$ is a $p$-modular system $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ such that $\mathcal{O}$ is a subring of $\mathcal{O}^{\prime}$ with $J(\mathcal{O})\subseteq J(\mathcal{O}^{\prime})$. For an algebra $A$, we denote by $A^{\rm op}$ the opposite algebra of $A$, by $A^{\times}$ the set of units in $A$, by $A$-mod the category of finitely generated $A$-modules, and by $K^{b}(A)$ the homotopy category of bound complexes over $A$-mod (see e.g. [23, page 117]). Unless otherwise stated, modules in this paper are left modules. By a block of $\mathcal{O}G$ (resp. $kG$) we usually mean an indecomposable direct factor $B$ of $\mathcal{O}G$ (resp. $kG$). Some times we also mean the correspondent primitive central idempotent.
Let $R\in\{\mathcal{O},k\}$ and $A$ a finitely generated $R$-algebra. A projective cover $(P,\varphi)$ of a finitely generated $A$-module $M$ is a projective $A$-module $P$ and a surjective $A$-module homomorphism $\varphi:P\to M$ such that no proper direct summand of $P$ surjects onto $M$. The assumption on $A$ ensure that projective covers of $M$ always exist and are unique up to isomorphism. The $A$-module $\Omega_{A}(M):={\rm ker}(\varphi)$ is called the Heller translate of $M$. We inductively define powers of the Heller translate $\Omega_{A}^{n}$, for a positive integer $n$. Finally, $\Omega_{A}^{0}(M)$ is defined as the direct sum of all the non-projective indecomposable summands of $M$.
@@
Superalgebras and supermodules. We review some definitions and notation from [22, §3]. Let $R$ be a commutative ring. An $R$-superspace is a finitely generated $R$-module $V$ with decomposition $V=V_{\bar{0}}\oplus V_{\bar{1}}$. For $\varepsilon\in\mathbb{Z}/2\mathbb{Z}$, we refer to the elements of $V_{\varepsilon}$ as the homogeneous elements of parity $\varepsilon$, and write $|v|=\varepsilon$ for $v\in V_{\varepsilon}$. The subspace $V_{\bar{0}}$ (resp. $V_{\bar{1}}$) is called the even (resp. odd ) part of $V$. An $R$-subsuperspace $W$ of $V$ is an $R$-submodule of $V$ such that $W=(W\cap V_{\bar{0}})\oplus(W\cap V_{\bar{1}})$. Of course we can treat $W$ as an $R$-subsuperspace in its own right. Let $V$ and $W$ be $R$-superspaces. The direct sum $V\oplus W$ is considered as an $R$-superspace in an obvious way. The tensor product $V\otimes_{R}W$ is considered as an $R$-superspace via $|v\otimes w|:=|v|+|w|$. A homomorphism $f:V\to W$ of $R$-superspaces is an $R$-linear map satisfying $f(V_{\varepsilon})\subseteq W_{\varepsilon}$, for $\varepsilon\in\mathbb{Z}/2\mathbb{Z}$.
An $R$-superalgebra is an $R$-superspace AA that is also an $R$-algebra such that $A_{\varepsilon}A_{\delta}\subseteq A_{\varepsilon+\delta}$ for any $\varepsilon,\delta\in\mathbb{Z}/2\mathbb{Z}$. A homomorphism $f:A\to B$ of $R$-superalgebras is an algebra homomorphism that is also a homomorphism of $R$-superspaces. Let $A$ be an $R$-superalgebra. If $A_{\bar{1}}\cap A^{\times}\neq\emptyset$, then $A$ is called a superalgebra with superunit, and any $u\in A_{\bar{1}}\cap A^{\times}$ is called a superunit.
An $A$-supermodule $M$ is an $A$-module which is also an $R$-superspace such that $A_{\varepsilon}M_{\delta}\subseteq M_{\varepsilon+\delta}$ for any $\varepsilon,\delta\in\mathbb{Z}/2\mathbb{Z}$. If $B$ is another $R$-superalgebra, an AA-$B$-bisupermodule module $M$ is an AA-$B$-bimodule which is also an $R$-superspace such that $A_{\varepsilon}M_{\delta}\subseteq M_{\varepsilon+\delta}$ and $M_{\delta}B_{\varepsilon}\subseteq M_{\varepsilon+\delta}$ for any $\varepsilon,\delta\in\mathbb{Z}/2\mathbb{Z}$. A homomorphism $f:M\to N$ of $A$-supermodules is an $A$-module homomorphism that is also a homomorphism of $R$-superspaces; a homomorphism of bisupermodules is defined similarly.
@@
Note that any finite generated algebra (module, bimodule) $A$ can be regarded as an superalgebra (a supermodule, a bisupermodule) with $A$ as the even part and 0 as the odd part. So the above tensor products, as well as most other results in this paper, for superalgebras, supermodules and bisupermodules apply to usual algebras, modules and bimodules.
@@
Morita superequivalences. Let $R$ be a commutative ring and let $A$, $B$ be $R$-superalgebras. Following [22], a Morita superequivalence between $A$ and $B$ is a Morita equivalence between $A$ and $B$ induced by an $(A,B)$-bisupermodule $M$ and a $(B,A)$-bisupermodule NN, i.e. $M\otimes_{B}N\cong A$ as $(A,A)$-bisupermodules and $N\otimes_{A}M\cong B$ as $(B,B)$-bisupermodules.
@@
We treat $M^{*}$ as a right $A$-supermodule via $(f\cdot a)(m):=f(am)$, for all $a\in A$, $f\in M^{*}$ and $m\in M$, hence also as an $A^{\rm sop}$-module via $(a\cdot f)(m):=(-1)^{|a||f|}f(am)$. For more properties of dual supermodules and relevant results on Morita superequivalences, we refer to [22, §3.7].
@@
If $A$ and $B$ are $G$-graded crossed $R$-superproducts, following [22, (3.5)] we define
@@
Let $R\in\{\mathcal{O},k\}$ and let $n$ be a positive integer. The twisted group superalgebra $\mathcal{T}_{n}^{R}$ over $R$ of the symmetric group $S_{n}$ is the $R$-superalgebra given by odd generators $\tau_{1},\cdots,\tau_{n-1}$ subject to the relations
@@
For each $w\in S_{n}$, choosing a reduced expression $w=s_{r_{1}}\cdots s_{r_{k}}$ in terms of simple transpositions, we define $\tau_{w}:=\tau_{r_{1}}\cdots\tau_{r_{k}}$ (which depends up to a sign on the choice of a reduced expression). It is well-known that $\{\tau_{w}\mid w\in S_{n}\}$ is an $R$-basis of $\mathcal{T}_{n}^{R}$ (one way to see that is to use the first isomorphism in (5.2) below).
Given an $R$-superalgebra AA, following [21, Section 5.1.a], we define the twisted wreath superproduct $A\wr_{s}\mathcal{T}_{n}^{R}$ as the free product $A^{\otimes n}\star\mathcal{T}_{n}^{R}$ of superalgebras subject to the relations
@@
$A\wr_{s}\mathcal{T}_{n}^{R}=\bigoplus_{w\in S_{n}}A^{\otimes n}\tau_{w},$ (2.3)
where $A^{\otimes n}$ denotes $\underbrace{A\otimes_{R}\cdots\otimes_{R}A}_{n}$. Hence $A\wr_{s}\mathcal{T}_{n}^{R}$ is an $S_{n}$-graded crossed $R$-superproduct.
Proof. Let $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be a $p$-modular system which is an extension of $(K,\mathcal{O},k)$ such that $\sqrt{-2}$ (a root of the polynomial $x^{2}+2\in\mathcal{O}[x]$) is contained in $\mathcal{O}^{\prime}$. Let $R^{\prime}\in\{\mathcal{O}^{\prime},k^{\prime}\}$ and let $A^{\prime}=R^{\prime}\otimes_{R}A$. Then by [22, (3.8)], $A^{\prime}\wr_{s}\mathcal{T}_{n}^{R^{\prime}}$ is an $S_{n}$-graded crossed $R^{\prime}$-superproduct via
$A^{\prime}\wr_{s}\mathcal{T}_{n}^{R^{\prime}}=\bigoplus_{w\in S_{n}}(A^{\prime})^{\otimes n}\tau_{w}.$ (2.4)
@@
To explain the equality (2.4), [22, (3.8)] uses the element $\sqrt{-2}$; see [22, Proposition 3.22]. But after the explanation, $\sqrt{-2}$ can be throw away - because $\sqrt{-2}$ would not involved in this equality.
@@
Keep the notation of Proposition 2.8. Given the $S_{n}$-grading in (2.3), recall the definition of $(A\wr_{s}\mathcal{T}_{n}^{R},A\wr_{s}\mathcal{T}_{n}^{R})_{S_{n}}$ from (2.2). Let $M$ be an $(A,A)$-bisupermodule. Following [22, (3.9)], we denote by $M^{\boxtimes n}_{S_{n}}$ the $(A\wr_{s}\mathcal{T}_{d}^{R},A\wr_{s}\mathcal{T}_{n}^{R})_{S_{n}}$-supermodule given by extending the $(A^{\otimes n},A^{\otimes n})$-bisupermodule $M^{\boxtimes n}:=\underbrace{M\boxtimes_{R}\cdots\boxtimes_{R}M}_{n}$, via
@@
Following [22, (3.10)] we now define
$M\wr_{s}\mathcal{T}_{n}^{R}:={\rm Ind}_{(A\wr_{s}\mathcal{T}_{n}^{R},A\wr_{s}\mathcal{T}_{n}^{R})_{S_{n}}}^{(A\wr_{s}\mathcal{T}_{n}^{R})\otimes_{R}(A\wr_{s}\mathcal{T}_{n}^{R})^{\rm sop}}(M^{\boxtimes n}_{S_{n}}).$ (2.6)
@@
Let $R\in\{\mathcal{O},k\}$, $n$ a positive integer and $A$ an $R$-superalgebra with a superunit $u$. If $C_{2}$ is a cyclic group with generator $g$, we can give $A$ the structure of a $C_{2}$-graded crossed $R$-superproduct with graded components $A_{g^{\varepsilon}}=A_{\bar{0}}u^{\varepsilon}=A_{\varepsilon}$ for $\varepsilon\in\mathbb{Z}/2\mathbb{Z}$. For any $n\in\mathbb{Z}_{>0}$, following [21, 5.1.c], we consider the wreath product
@@
where $g_{r}$ is the generator of the $r$-th $C_{2}$ factor in the base group $C_{2}^{\times n}$. Then we can give $A\wr_{s}\mathcal{T}_{n}^{R}$ the structure of a $(C_{2}\wr S_{n})$-graded crossed $R$-superproduct with graded components
$(A\wr_{s}\mathcal{T}_{n})_{g_{1}^{\varepsilon_{1}}\cdots g_{n}^{\varepsilon_{n}}w}=A_{\bar{0}}^{\otimes n}u_{1}^{\varepsilon_{1}}\dots u_{n}^{\varepsilon_{n}}\tau_{w}$
@@
$(A\wr_{s}\mathcal{T}_{n}^{R},B\wr_{s}\mathcal{T}_{n}^{R})_{C_{2}\wr S_{n}}\subseteq(A\wr_{s}\mathcal{T}_{n}^{R})\otimes_{R}(B\wr_{s}\mathcal{T}_{n}^{R})^{\rm sop}.$
Setting $U_{r}:=(u_{A})_{r}\otimes(u_{B}^{-1})_{r}$, for $1\leq r\leq n$, we have that $A_{\bar{0}}^{\otimes n}\otimes_{R}(B_{\bar{0}}^{\otimes n})^{\rm sop}$ together with the $U_{r}$’s generate a subsuperalgebra of $(A\wr_{s}\mathcal{T}_{n}^{R},B\wr_{s}\mathcal{T}_{n}^{R})_{C_{2}\wr_{s}S_{n}}$ isomorphic to $((A,B)_{C_{2}})^{\otimes n}$. We identify the two superalgebras. Then $(A\wr_{s}\mathcal{T}_{n}^{R},B\wr_{s}\mathcal{T}_{n}^{R})_{C_{2}\wr_{s}S_{n}}$ is generated by $((A,B)_{C_{2}})^{\otimes n}$ and $T_{r}:=\tau_{r}\otimes\tau_{r}^{-1}$, for $1\leq r\leq n-1$ subject to the relations listed in [21, (5.1.13)].
Proposition 2.12 (a slight generalisation of [21, Proposition 5.1.14]).
Let $R\in\{\mathcal{O},k\}$, let $n$ be a positive integer, and let $A$ and $B$ be $R$-superalgebras with superunit. If $X$ is a complex of $A_{\bar{0}}$-$B_{\bar{0}}$-bimodules that extends to a complex of $(A,B)_{C_{2}}$-modules, then $X^{\otimes n}$ extends to a complex of $(A\wr_{s}\mathcal{T}_{n},B\wr_{s}\mathcal{T}_{n})_{C_{2}\wr S_{n}}$-modules.

### v5

@@ 1 Introduction
Let $0<d<p$ and $\rho$ be a $d$-Rouquier $\bar{p}$-core. Then $B^{\eta,\rho,d}_{R_{p}}$ and $(B^{\eta,\rho,d}_{R_{p}})_{\bar{0}}$ are splendidly Rickard equivalent to $b^{\eta,\rho,d}_{R_{p}}$ and $(b^{\eta,\rho,d}_{R_{p}})_{\bar{0}}$ respectively. In particular, Conjecture 1.1 holds for RoCK blocks of $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$.
@@
Using [1, Theorem 1.4] and [15, Theorem 1], we obtain a corollary of Theorem 1.9:
@@
Refinements of the Alperin–McKay conjecture by Isaacs–Navarro, Navarro, Turull and Boltje (see [1, (1.5), (1.6), (1.7), (1.8)]) and the blockwise Navarro/Galois Alperin weight conjecture (see [15, Conjecture 2]) hold for RoCK blocks of $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$.
The blockwise Navarro Alperin weight conjecture has been proved for all blocks of $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$ in [9] using different method.
For a partition $\lambda$ we denote by $h(\lambda)$ the number of nonzero parts of $\lambda$. Now by Theorem 1.9 and Theorem 5.10 below, to prove Conjecture 1.1 for all spin blocks of symmetric and alternating groups, it suffices to prove the following refined version of the Kessar–Schaps conjecture [21]:
@@
Arbitrary spin blocks $B^{\eta,\rho,d}_{R_{p}}$ and $B^{\eta,\rho^{\prime},d^{\prime}}_{R_{p}}$ of symmetric groups (or spin blocks $(B^{\eta,\rho,d}_{R_{p}})_{\bar{0}}$ and $(B^{\eta,\rho^{\prime},d^{\prime}}_{R_{p}})_{\bar{0}}$ of alternating groups) are splendidly Rickard equivalent if and only if $d=d^{\prime}$ and $|\rho|-h(\rho)\equiv|\rho^{\prime}|-h(\rho^{\prime})~({\rm mod}~2)$.
@@
This paper is organised as follows: In Section 2 we review some relevant results on superalgebras and supermodules which are necessary for later use. In Section 3 we prove descent results for twisted group superalgebras and twisted wreath superproducts. In Section 4, we prove general descent results for supermodules and superequivalences. In Section 5, we formulate convenient reference for some useful results on graded Rickard equivalences and splendid complexes. In Section 6 we review some block theory of $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$. Based on [22], we prove in Section 7 the structure results for Brauer correspondents of spin blocks over $\mathbb{F}_{p}$. In Section 8, we analyse the Morita superequivalence bimodule for an RoCK block constructed by Kleshchev and Livesey [23], and then realise this bisupermodule over $\mathbb{Z}_{p}$; we prove the refined Broué conjecture for RoCK spin blocks in the end of Section 8.
@@ 2 Preliminaries
Generalities. Throughout the rest of this paper $p$ is an odd prime, $\mathcal{O}$ is a complete discrete valuation ring with residue field $k$ of characteristic $p$ and quotient field $K$ of characteristic 0. Such a triple $(K,\mathcal{O},k)$ is called a $p$-modular system. (Note that we didn’t assume $(K,\mathcal{O},k)$ to be large enough.) An extension of the $p$-modular system $(K,\mathcal{O},k)$ is a $p$-modular system $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ such that $\mathcal{O}$ is a subring of $\mathcal{O}^{\prime}$ with $J(\mathcal{O})\subseteq J(\mathcal{O}^{\prime})$. For an algebra $A$, we denote by $A^{\rm op}$ the opposite algebra of $A$, by $A^{\times}$ the set of units in $A$, by $A$-mod the category of finitely generated $A$-modules, and by $K^{b}(A)$ the homotopy category of bound complexes over $A$-mod (see e.g. [24, page 117]). Unless otherwise stated, modules in this paper are left modules. By a block of $\mathcal{O}G$ (resp. $kG$) we usually mean an indecomposable direct factor $B$ of $\mathcal{O}G$ (resp. $kG$); some times we also mean the corresponding primitive central idempotent.
Let $R\in\{\mathcal{O},k\}$ and $A$ a finitely generated $R$-algebra. A projective cover $(P,\varphi)$ of a finitely generated $A$-module $M$ is a projective $A$-module $P$ and a surjective $A$-module homomorphism $\varphi:P\to M$ such that no proper direct summand of $P$ surjects onto $M$. The assumption on $A$ ensures that projective covers of $M$ always exist and are unique up to isomorphism. The $A$-module $\Omega_{A}(M):={\rm ker}(\varphi)$ is called the Heller translate of $M$. We inductively define powers of the Heller translate $\Omega_{A}^{n}$, for a positive integer $n$. Finally, $\Omega_{A}^{0}(M)$ is defined as the direct sum of all the non-projective indecomposable summands of $M$.
@@
Superalgebras and supermodules. We review some definitions and notation from [23, §3]. Let $R$ be a commutative ring. An $R$-superspace is a finitely generated $R$-module $V$ with decomposition $V=V_{\bar{0}}\oplus V_{\bar{1}}$. For $\varepsilon\in\mathbb{Z}/2\mathbb{Z}$, we refer to the elements of $V_{\varepsilon}$ as the homogeneous elements of parity $\varepsilon$, and write $|v|=\varepsilon$ for $v\in V_{\varepsilon}$. The subspace $V_{\bar{0}}$ (resp. $V_{\bar{1}}$) is called the even (resp. odd ) part of $V$. An $R$-subsuperspace $W$ of $V$ is an $R$-submodule of $V$ such that $W=(W\cap V_{\bar{0}})\oplus(W\cap V_{\bar{1}})$. Of course we can treat $W$ as an $R$-subsuperspace in its own right. Let $V$ and $W$ be $R$-superspaces. The direct sum $V\oplus W$ is considered as an $R$-superspace in an obvious way. The tensor product $V\otimes_{R}W$ is considered as an $R$-superspace via $|v\otimes w|:=|v|+|w|$. A homomorphism $f:V\to W$ of $R$-superspaces is an $R$-linear map satisfying $f(V_{\varepsilon})\subseteq W_{\varepsilon}$, for $\varepsilon\in\mathbb{Z}/2\mathbb{Z}$.
An $R$-superalgebra is an $R$-superspace AA that is also an $R$-algebra such that $A_{\varepsilon}A_{\delta}\subseteq A_{\varepsilon+\delta}$ for any $\varepsilon,\delta\in\mathbb{Z}/2\mathbb{Z}$. A homomorphism $f:A\to B$ of $R$-superalgebras is an algebra homomorphism that is also a homomorphism of $R$-superspaces. Let $A$ be an $R$-superalgebra. If $A_{\bar{1}}\cap A^{\times}\neq\varnothing$, then $A$ is called a superalgebra with superunit, and any $u\in A_{\bar{1}}\cap A^{\times}$ is called a superunit.
An $A$-supermodule $M$ is an $A$-module which is also an $R$-superspace such that $A_{\varepsilon}M_{\delta}\subseteq M_{\varepsilon+\delta}$ for any $\varepsilon,\delta\in\mathbb{Z}/2\mathbb{Z}$. If $B$ is another $R$-superalgebra, an AA-$B$-bisupermodule module $M$ is an AA-$B$-bimodule which is also an $R$-superspace such that $A_{\varepsilon}M_{\delta}\subseteq M_{\varepsilon+\delta}$ and $M_{\delta}B_{\varepsilon}\subseteq M_{\delta+\varepsilon}$ for any $\varepsilon,\delta\in\mathbb{Z}/2\mathbb{Z}$. A homomorphism $f:M\to N$ of $A$-supermodules is an $A$-module homomorphism that is also a homomorphism of $R$-superspaces; a homomorphism of bisupermodules is defined similarly.
@@
Note that any finitely generated algebra (module, bimodule) $A$ can be regarded as a superalgebra (supermodule, bisupermodule) with $A$ as the even part and 0 as the odd part. So the above tensor products (as well as most other results in this paper) for superalgebras, supermodules and bisupermodules apply to usual algebras, modules and bimodules.
@@
Morita superequivalences. Let $R$ be a commutative ring and let $A$, $B$ be $R$-superalgebras. Following [23], a Morita superequivalence between $A$ and $B$ is a Morita equivalence between $A$ and $B$ induced by an $(A,B)$-bisupermodule $M$ and a $(B,A)$-bisupermodule NN, i.e. $M\otimes_{B}N\cong A$ as $(A,A)$-bisupermodules and $N\otimes_{A}M\cong B$ as $(B,B)$-bisupermodules. Here $M\otimes_{B}N$ is an $(A,A)$-bisupermodule with $(M\otimes_{B}N)_{\bar{0}}=M_{\bar{0}}\otimes_{B}N_{\bar{0}}\oplus M_{\bar{1}}\otimes_{B}N_{\bar{1}}$ and $(M\otimes_{B}N)_{\bar{1}}=M_{\bar{0}}\otimes_{B}N_{\bar{1}}\oplus M_{\bar{1}}\otimes_{B}N_{\bar{0}}$.
@@
We treat $M^{*}$ as a right $A$-supermodule via $(f\cdot a)(m):=f(am)$, for all $a\in A$, $f\in M^{*}$ and $m\in M$, hence also as an $A^{\rm sop}$-module via $(a\cdot f)(m):=(-1)^{|a||f|}f(am)$. For more properties of dual supermodules and relevant results on Morita superequivalences, we refer to [23, §3.7].
@@
If $A$ and $B$ are $G$-graded crossed $R$-superproducts, following [23, (3.5)] we define
@@
Let $R\in\{\mathcal{O},k\}$ and let $n$ be a positive integer. In [22, 23], Kleshchev and Livesey defined the twisted group superalgebra $\mathcal{T}_{n,R}$ over $R$ of the symmetric group $S_{n}$ to be the $R$-superalgebra given by odd generators $\tau_{1},\cdots,\tau_{n-1}$ subject to the relations
@@
For each $w\in S_{n}$, choosing a reduced expression $w=s_{r_{1}}\cdots s_{r_{k}}$ in terms of simple transpositions, we define $\tau_{w}:=\tau_{r_{1}}\cdots\tau_{r_{k}}$ (which depends up to a sign on the choice of a reduced expression). It is well-known that $\{\tau_{w}\mid w\in S_{n}\}$ is an $R$-basis of $\mathcal{T}_{n,R}$ (one way to see that is to use the first isomorphism in (6.2) below).
Given an $R$-superalgebra AA, Kleshchev and Livesey [22, Section 5.1.a] defined the twisted wreath superproduct $A\wr_{s}\mathcal{T}_{n,R}$ as the free product $A^{\otimes n}\star\mathcal{T}_{n,R}$ of superalgebras subject to the relations
@@
$A\wr_{s}\mathcal{T}_{n,R}=\bigoplus_{w\in S_{n}}A^{\otimes n}\tau_{w},$ (2.3)
where $A^{\otimes n}$ denotes $\underbrace{A\otimes_{R}\cdots\otimes_{R}A}_{n}$. Hence $A\wr_{s}\mathcal{T}_{n,R}$ is an $S_{n}$-graded crossed $R$-superproduct.
Proof. Let $(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be a $p$-modular system which is an extension of $(K,\mathcal{O},k)$ such that $\sqrt{-2}$ (a root of the polynomial $x^{2}+2\in\mathcal{O}[x]$) is contained in $\mathcal{O}^{\prime}$. Let $R^{\prime}\in\{\mathcal{O}^{\prime},k^{\prime}\}$ and let $A^{\prime}=R^{\prime}\otimes_{R}A$. Then by [23, (3.8)], $A^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}}$ is an $S_{n}$-graded crossed $R^{\prime}$-superproduct via
$A^{\prime}\wr_{s}\mathcal{T}_{n,R^{\prime}}=\bigoplus_{w\in S_{n}}(A^{\prime})^{\otimes n}\tau_{w}.$ (2.4)
@@
To explain the equality (2.4), [23, (3.8)] uses the element $\sqrt{-2}$; see [23, Proposition 3.22]. But after the explanation, $\sqrt{-2}$ can be throw away - because $\sqrt{-2}$ would not involved in this equality.
@@
Keep the notation of Proposition 2.8. Given the $S_{n}$-grading in (2.3), recall the definition of $(A\wr_{s}\mathcal{T}_{n,R},A\wr_{s}\mathcal{T}_{n,R})_{S_{n}}$ from (2.2). Let $M$ be an $(A,A)$-bisupermodule. Following [23, (3.9)], we denote by $M^{\boxtimes n}_{S_{n}}$ the $(A\wr_{s}\mathcal{T}_{n,R},A\wr_{s}\mathcal{T}_{n,R})_{S_{n}}$-supermodule given by extending the $(A^{\otimes n},A^{\otimes n})$-bisupermodule $M^{\boxtimes n}:=\underbrace{M\boxtimes_{R}\cdots\boxtimes_{R}M}_{n}$, via
@@
Following [23, (3.10)] we now define
$M\wr_{s}\mathcal{T}_{n,R}:={\rm Ind}_{(A\wr_{s}\mathcal{T}_{n,R},A\wr_{s}\mathcal{T}_{n,R})_{S_{n}}}^{(A\wr_{s}\mathcal{T}_{n,R})\otimes_{R}(A\wr_{s}\mathcal{T}_{n,R})^{\rm sop}}(M^{\boxtimes n}_{S_{n}}).$ (2.6)
@@
Let $R\in\{\mathcal{O},k\}$, $n$ a positive integer and $A$ an $R$-superalgebra with a superunit $u$. If $C_{2}$ is a cyclic group of order $2$ with generator $g$, we can give $A$ the structure of a $C_{2}$-graded crossed $R$-superproduct with graded components $A_{g^{\varepsilon}}=A_{\bar{0}}u^{\varepsilon}=A_{\varepsilon}$ for $\varepsilon\in\mathbb{Z}/2\mathbb{Z}$. For any $n\in\mathbb{Z}_{>0}$, following [22, 5.1.c], we consider the wreath product
@@
where $g_{r}$ is the generator of the $r$-th $C_{2}$ factor in the base group $C_{2}^{\times n}$. Then we can give $A\wr_{s}\mathcal{T}_{n,R}$ the structure of a crossed product of $A_{\bar{0}}^{\otimes n}$ and $C_{2}\wr S_{n}$ (see 5.1) with graded components
$(A\wr_{s}\mathcal{T}_{n,R})_{g_{1}^{\varepsilon_{1}}\cdots g_{n}^{\varepsilon_{n}}w}=A_{\bar{0}}^{\otimes n}u_{1}^{\varepsilon_{1}}\dots u_{n}^{\varepsilon_{n}}\tau_{w}$
@@
$(A\wr_{s}\mathcal{T}_{n,R},B\wr_{s}\mathcal{T}_{n,R})_{C_{2}\wr S_{n}}\subseteq(A\wr_{s}\mathcal{T}_{n,R})\otimes_{R}(B\wr_{s}\mathcal{T}_{n,R})^{\rm sop}.$
Setting $U_{r}:=(u_{A})_{r}\otimes(u_{B}^{-1})_{r}$, for $1\leq r\leq n$, we have that $A_{\bar{0}}^{\otimes n}\otimes_{R}(B_{\bar{0}}^{\otimes n})^{\rm sop}$ together with the $U_{r}$’s generate a subsuperalgebra of $(A\wr_{s}\mathcal{T}_{n,R},B\wr_{s}\mathcal{T}_{n,R})_{C_{2}\wr_{s}S_{n}}$ isomorphic to $((A,B)_{C_{2}})^{\otimes n}$. We identify the two superalgebras. Then $(A\wr_{s}\mathcal{T}_{n,R},B\wr_{s}\mathcal{T}_{n,R})_{C_{2}\wr_{s}S_{n}}$ is generated by $((A,B)_{C_{2}})^{\otimes n}$ and $\{T_{r}:=\tau_{r}\otimes\tau_{r}^{-1}\mid 1\leq r\leq n-1\}$ subject to the relations listed in [22, (5.1.13)].
Proposition 2.12 (a slight generalisation of [22, Proposition 5.1.14]).
Let $R\in\{\mathcal{O},k\}$, let $n$ be a positive integer, and let $A$ and $B$ be $R$-superalgebras with superunit. If $X$ is a complex of $A_{\bar{0}}$-$B_{\bar{0}}$-bimodules that extends to a complex of $(A,B)_{C_{2}}$-modules, then $X^{\otimes n}$ extends to a complex of $(A\wr_{s}\mathcal{T}_{n,R},B\wr_{s}\mathcal{T}_{n,R})_{C_{2}\wr S_{n}}$-modules.